OpenStudy (anonymous):

write each of the following in the form Ax+By=C 1.y=-x+5 2.x/4+y/5=1 3.10=9x-5y/3 please answer it .. in math equation..!

OpenStudy (anonymous):

i need a good answer

OpenStudy (anonymous):

1.x+y=5 2.5x+4y=20 3.27x-5y=30

OpenStudy (anonymous):

what is this????

OpenStudy (anonymous):

answers

OpenStudy (anonymous):

http://www.twiddla.com/583948 come to twiddla i will explain it to you

OpenStudy (anonymous):

this is wrong..

OpenStudy (anonymous):

this is right....

OpenStudy (anonymous):

and you cant prove me wrong ........ thats a deal

OpenStudy (anonymous):

show me your equations..

OpenStudy (anonymous):

1.x+y=5 2.5x+4y=20 3.27x-5y=30